OpenStudy (anonymous):

determine whether lim exists for f(x,y) = x^4/(x^2+y^2) as (x,y)-> (0,0)

OpenStudy (anonymous):

Have you tried converting to polar?

OpenStudy (anonymous):

no i haven't is it necessary ?

OpenStudy (anonymous):

Might not be, but that's my first impression. Converting the polar changes the denominator to \(r^2\) and the numerator to \(r^4\cos^4\theta\). Then as \((x,y)\to(0,0)\), I believe it would be \((r,\theta)\to(0,0)\).

OpenStudy (anonymous):

polars i tend to forget but i see what you mean,

OpenStudy (anonymous):

Wolfram says the limit turns out to be 0, which works well with my suggestion.

OpenStudy (anonymous):

i did it too on wolfram jus wanted more depth into how, can also you replace y=x

OpenStudy (anonymous):

making it into x^4/2x^2 giving you zero

OpenStudy (anonymous):

Yes, you could do that, but there's no guarantee that the result of that will give you the right answer. You have to consider "all" possible paths as \((x,y)\to(0,0)\). For example (not regarding this limit, but in general), substituting \(y=x\) may give you 0, but substituting \(x=0\) might not, and so on. Which is why I prefer the polar conversion method.

OpenStudy (anonymous):

okay thanks ill keep that in mind

OpenStudy (anonymous):

You're welcome.